# Splat-MOVER: Multi-Stage, Open-Vocabulary Robotic Manipulation via Editable Gaussian Splatting

### v2

@@ V-A Semantic Localization via ASK-Splat
Given a natural-language query specifying an object of interest, SEE-Splat leverages ASK-Splat to identify semantically relevant Gaussians, as discussed in Section IV-A. Figure 2 shows the localization of an electric stove, saucepan, and a fruit in a real-world Kitchen scene. To improve the localization accuracy, the text prompt can include the geometric and visual properties of the object, such as its color, in addition to its semantic class. At this stage, SEE-Splat generates a semantic similarity map, from which relevant Gaussians are extracted, given a threshold on the semantic score.
V-B Masking the Gaussians in SEE-Splat
@@
Given a semantic similarity map of the scene, SEE-Splat generates a mask identifying the Gaussians relevant to the specified object. This procedure begins with thresholding the semantic scores of each Gaussian to remove dissimilar Gaussians from the set of relevant Gaussians, which creates a sparse point cloud of the relevant Gaussians, constructed from the means of these Gaussians. However, photo-realistic rendering of Gaussian environments require denser point clouds. Consequently, SEE-Splat lifts the features of the point cloud from the $3$D Euclidean space to a $7$D feature space, by augmenting each point in the point cloud with its RGB color and semantic score. Subsequently, SEE-Splat identifies all neighboring points in the scene within a specified distance of the point cloud in the $7$D feature space using an efficient KD-tree query. SEE-Splat incorporates these points into the point cloud to create a denser point cloud, comprising of all semantically-relevant Gaussians, while removing outliers from the set of points. In Figure 2, we show the Gaussians extracted by SEE-Splat, as a point cloud with well-defined geometry, given a natural-language query for each object.
V-C Editing the Gaussians in SEE-Splat
@@
Leveraging the Gaussian primitives in ASK-Splat, SEE-Splat enables real-time scene-editing by inserting new Gaussians into the scene, removing Gaussians, and modifying the properties of the Gaussians, to reflect (or simulate) changes in the real-world. SEE-Splat supports seamless insertion and removal of Gaussians by introducing or deleting the relevant Gaussians from the set of Gaussians representing the scene, respectively. In addition, SEE-Splat supports both rigid and non-rigid transformation of the Gaussians, enabling simulated motion of the Gaussians, as well as changes to the shape of the Gaussians via non-isometric scaling. Specifically, given a function specifying the transformation ${\xi:\mathcal{G}_{s}\mapsto\mathcal{G}_{s}}$ (where $\mathcal{G}_{s}$ represents the space of the Gaussian primitives), SEE-Splat updates the spatial attributes of the relevant Gaussians by applying $\xi$ to these Gaussians. In the case of rigid transformations, ${\xi}$ can be described by an $\mathrm{SE}(3)$ transformation matrix, specifying rotation and translation of the Gaussians. We can render the edited scene to generate photo-realistic visualizations. Although, we do not consider physics-based simulations in this work, we note that physics can be incorporated into SEE-Splat to achieve realism. Deletion and transformation of the Gaussians introduces artifacts into the scene representation, degrading its photo-realistic qualities. To address this challenge, SEE-Splat enables $3$D Gaussian infilling by introducing new Gaussians with similar attributes in regions with missing geometry. Figure 3 provides an illustration of such artifacts (e.g., the hole in the table), when the scene is edited to visualize the effects of moving the saucepan to the electric stove. Through $3$D Gaussian infilling, SEE-Splat generates a photorealistic rendering of the edited scene, eliminating these artifacts.
@@
VI Grasping and Manipulation with Splat-MOVER
Via Splat-MOVER, we provide a concrete application of ASK-Splat and SEE-Splat to multi-stage robotic manipulation problems. We consider language-guided, multi-stage robotic grasping and placing tasks, where each stage consists of grasping and placing distinct objects within a scene, which arise, for example, in scenarios where robotic manipulators are utilized in meal preparation or cleanup. Specifically, given a language description of a task, we seek to design an algorithm that endows a robotic manipulator with spatial and semantic understanding of its environment to enable it to complete a specified grasping and placing task, while incorporating the temporal effects of its actions in its representation of its environment to facilitate the completion of downstream stages. We assume that the manipulator is equipped with a parallel-jaw gripper; however, we note that the algorithm in this work can be extended to other gripper forms. Building upon ASK-Splat and SEE-Splat, Splat-MOVER introduces Grasp-Splat, a grasp generation and selection module for robotic manipulation, described in the following discussion.
VI-A Grasp-Splat for Grasp Proposal
Grasp-Splat utilizes the dense point cloud of the object generated by SEE-Splat (e.g., Figure 2) to propose grasp configurations for grasping the object, while harnessing the grasp affordance knowledge in ASK-Splat to identify more compatible grasp configurations associated with the specified object. The proposed grasps are generated from the point cloud using GraspNet [8], a learning-based approach that utilizes an encoder-decoder architecture to estimate $6D$ grasp poses for a parallel-jaw gripper from a point cloud of the object, along with estimated grasp-quality scores associated with these grasp poses.
@@ VI-A Grasp-Splat for Grasp Proposal
We note that the grasps generated by GraspNet are not always ideal. For example, the grasps generated by GraspNet in Figure 4 are either infeasible or challenging to execute. Consequently, Grasp-Splat executes a grasp selection procedure to identify more-promising candidate grasps. We hypothesize that leveraging grasp affordance of each part of the object when generating candidate grasps could be essential to identifying better candidate grasps. As a result, we introduce the grasp metric ${\nu:\mathrm{SE}(3)\mapsto\mathbb{R}}$, which computes the grasp affordance at a specified grasp pose ${X\in\mathrm{SE}(3)}$. Grasp-Splat ranks the candidate grasps generated by GraspNet based on the grasp scores given by $\nu(X)$, leveraging the affordance associated with each grasp pose to identify grasp configurations that are more likely to succeed, depicted in Figure 4. Moreover, since GraspNet does not consider the position of the robot relative to the object, the proposed grasps might require post-processing to account for the relative position of the robot.
VI-B Multi-Stage Robotic Manipulation
For multi-stage robotic manipulation, we begin by decomposing the manipulation task into stages. Our approach supports the specification of the stages comprising the task by a human or by a large language model (LLM). In the case where the natural-language description of the task does not specify the stages involved in the task, we query an LLM for the stages required to complete the manipulation task. For each stage in the manipulation task, we utilize ASK-Splat, SEE-Splat, and Grasp-Splat to identify the relevant object and generate candidate grasp poses, as described in Sections IV and VI-A. Likewise, we query ASK-Splat for the target location for placing the object. We evaluate the feasibility of each candidate grasp using an off-the-shelf motion planner for the robotic manipulator, inputting the point cloud of the scene, extracted from ASK-Splat and SEE-Splat, into the motion planner, which the motion planner uses for collision detection during motion planning. We execute the top candidate grasps, moving on to the next if the robot motion planner fails to compute a solution to execute the selected grasp.
@@ VI-B Multi-Stage Robotic Manipulation
VII Evaluations
We examine the effectiveness of ASK-Splat, SEE-Splat, and Splat-MOVER in open-vocabulary, multi-stage robotic manipulation problems. We begin with evaluating the affordance and semantic knowledge embedded in ASK-Splat, assessing the quality of the encoded grasp affordance and semantic masks generated by ASK-Splat across a variety of environments and objects.
VII-A Experimental Setup
VII-A1 ASK-Splat
@@ VII-A3 Splat-MOVER
Specifically, we evaluate Splat-MOVER in four multi-stage manipulation tasks across three scenes: the Kitchen, Cleaning, and Workshop scenes. In the Kitchen scene, we consider a Cooking task where the robot is asked to place a saucepan on an electric burner (Stage 1) and subsequently place a fruit inside the saucepan (Stage 2). Further, in the Kitchen scene, we consider a Chopping task where the robot is asked to place a knife on a chopping board (Stage 1) and subsequently place a fruit next to the knife (Stage 2). We consider a Cleaning task (in the Cleaning scene), where the robot is asked to place a cleaning spray in a bin (Stage 1) and subsequently place a sponge next to the cleaning spray (Stage 2). Lastly, in the Workshop scene, the robot is asked to place a power drill on a work mat (Stage 1) and subsequently place a wooden block next to the drill (Stage 2), which we refer to as the Workshop task. We show novel views of the scene, rendered from ASK-Splat, associated with each task in Figure 5.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Multi-stage Robotic Manipulation: (top-left) Cooking task, where a robot is asked to move a saucepan to an electric burner, followed by placing a fruit in the saucepan; (top-right) Chopping task, where a robot is asked to move a knife to a chopping board, followed by placing a fruit next to the knife; (bottom-left) Cleaning task, where a robot is asked to move a cleaning spray into a bin, followed by placing a sponge next to the bin; (bottom-right) Workshop task, where a robot is asked to move a power drill to a work mat, followed by moving a wooden block next to the drill.
VII-A4 Hardware Experiments
@@ VII-B ASK-Splat Representation
VII-C Splat-MOVER for Multi-Stage Robotic Manipulation
@@
We compare Splat-MOVER to the existing open-vocabulary robotic manipulation methods LERF-TOGO [9] and F3RM [5] in four tasks: the Cooking task, Chopping task, Cleaning task, and Workshop task, described in Section VII-A3. We utilize the publicly-available implementation of LERF-TOGO provided by the authors. Since F3RM requires collecting human demonstrations in virtual-reality, which we could not provide in our evaluations, we implement F3RM and utilize GraspNet for grasp generation. Since the NeRF-based methods LERF-TOGO and F3RM are not amenable to scene-editing, we limit our comparisons to the first stage of the manipulation task, where we consider grasping a saucepan, knife in a knife guard, cleaning spray, and power drill. Figure 8 shows a few candidate grasps proposed by GraspNet, F3RM, LERF-TOGO, and Grasp-Splat for each of these objects. GraspNet does not consider the semantic features of the object in generating candidate grasps; as a result, the proposed grasps are not localized in regions where a human might grasp the object, unlike the candidate grasps proposed by F3RM, LERF-TOGO, and Grasp-Splat, which generate grasps closer to the handle of the respective objects. For example, the proposed grasps lie relatively close to the handle of the saucepan and the knife. F3RM and LERF-TOGO generate candidate grasp conditioned on a text prompt identifying the region to grasp the object (such as its handle) provided by a human operator or an LLM (in LERF-TOGO) or from a dataset of human demonstrations (in F3RM). In contrast, Grasp-Splat does not require any external guidance to generate candidate grasps of similar quality, harnessing the grasp affordances provided by ASK-Splat. We summarize the capabilities of each of these methods in Table I.
TABLE I: Representation Capabilities of LERF-TOGO [9], F3RM [5], and Splat-MOVER
Capabilities Semantic Knowledge Affordance Knowledge Scene-Editing LERF-TOGO [9] F3RM [5] Splat-MOVER (ours)
In the hardware experiments on the robot, we describe a candidate grasp as being feasible if the MoveIt planner successfully finds a plan to execute the grasp. We execute $20$ feasible trials for each object, and in Table II, provide the grasping success rate of each method, as well as the percentage of grasps that lie within the affordance region of the object (AGSR), where the affordance region of each object is defined as follows: the entire handle of the saucepan, the entire knife with the guard on, the region of the cleaning spray excluding its cap, and the middle region of the power drill below its top compartment and above its battery compartment. From Table II, Splat-MOVER achieves the best grasping success rates and the highest percentage of grasps in the affordance region of each object in grasping the saucepan, knife, and cleaning spray. LERF-TOGO attains a perfect success rate in grasping the power drill; however, LERF-TOGO always grasped the top of the drill, outside its affordance region, depicted in Figure 8. In contrast, Splat-MOVER achieves a $95\%$ success rate, grasping the object within its affordance region.
In addition, we evaluate the pick-and-place success rate of all the methods in each manipulation task, where the place success rate is conditioned on the number of successful trials in picking the object. The publicly-available implementation of LERF-TOGO does not support the specification of a place location. Hence, we do not evaluate its place success rate. In Table III, we present the pick-and-place success rates in the Cooking task. Splat-MOVER achieves a perfect success rate in picking up the saucepan, with a place success rate of $60\%$. We note that the place step involves placing the saucepan on an electric burner, which is quite challenging, potentially explaining the lower place success rate. Nonetheless, Splat-MOVER outperforms LERF-TOGO and F3RM in the first stage of the task, the only stage of the task that both methods can perform.
Table IV provides the pick-and-place success rates in the Chopping task. Again, Splat-MOVER achieves the highest pick success rate ($85\%$) in Stage $1$ of the task. Although F3RM achieves the highest place success rate, it achieves a much lower pick success rate compared to Splat-MOVER. In addition, in the Cleaning and Workshop tasks, Splat-MOVER achieves the highest success rates in the first stage of each task, and further achieves relatively high success rates in the second stage of each task, shown in Tables V and VI. LERF-TOGO achieves a perfect success rate in picking up the power drill in the first stage of the Workshop task. Since LERF-TOGO and F3RM are not amenable to multi-stage manipulation tasks, we cannot evaluate the success rate of these methods for the entire manipulation task. In contrast, Splat-MOVER enables multi-stage robotic manipulation, achieving a task success rate of $40\%$, $65\%$, $70\%$, and $80\%$ in the Cooking, Chopping, Cleaning, and Workshop tasks, respectively. We note that the Cooking task is the most challenging task, compared to the other tasks, given the little margin of error tolerated in placing the saucepan on the electric burner.
@@
TABLE II: Grasping success rate and the percentage of grasps in the affordance region (AGSR) of each object for LERF-TOGO [9], F3RM [5], and Splat-MOVER across $20$ feasible trials. The top-performing stat is shown in bold.
[…]
TABLE III: Pick and Place Success Rates in a two-stage manipulation Cooking task, where the robot must move a saucepan to an electric burner (Stage 1), then move a fruit into the saucepan (Stage 2).
[…]
TABLE IV: Pick and Place Success Rates in a two-stage manipulation Chopping task, where the robot must move a knife to a chopping board (Stage 1), then move a fruit next to the knife (Stage 2).
[…]
TABLE V: Pick and Place Success Rates in a two-stage manipulation Cleaning task, where the robot must move a cleaning spray into a bin (Stage 1), then move a sponge next to the cleaning spray inside the bin (Stage 2).
[…]
TABLE VI: Pick and Place Success Rates in a two-stage manipulation in a Workshop task, where the robot is tasked with moving a power drill onto a work mat (Stage 1), followed by moving a wooden block next to the drill on the work mat (Stage 2).
[…]
VIII Conclusion
We present Splat-MOVER, a robotics stack for multi-stage open-vocabulary robotic manipulation. Splat-MOVER consists of three modules: ASK-Splat, SEE-Splat, and Grasp-Splat. ASK-Splat enables semantic and affordance queries via natural-language interaction to identify relevant objects within $3$D scenes, while SEE-Splat provides real-time, dynamic scene editing, enabling visualization of the evolution of the real-world scene due to the robot’s interaction with objects within the scene. Grasp-Splat builds upon these two modules for affordance-aware grasp generation, necessary for effective multi-stage robotic manipulation. We demonstrate the effectiveness of Splat-MOVER in real-world experiments in comparison to two recent baseline methods. Compared to the existing works, Splat-MOVER endows robots with the unique capability for multi-stage, open-vocabulary manipulation with minimal human inputs, by leveraging distilled grasp affordance knowledge and real-time dynamic scene editing, which are essential to multi-stage manipulation.
IX Limitations and Future Work
We present a few limitations of Splat-MOVER and provide directions for future work. ASK-Splat distills grasp affordance knowledge from foundation models into $3$D Gaussian Splatting scenes. We note that the effectiveness of the distilled grasp affordances and the generalization capability of our model is highly dependent on that of the affordance foundation model. For example, when trained using the VRB foundation model, an ASK-Splat scene may not generalize notably well to markedly-different objects outside of those seen in the EPIC-KITCHENS dataset. Future work will examine enhancing the generalization capability of the proposed ASK-Splat module by training and employing diverse cross-environment, cross-task, vision-affordance foundation models, rather than relying on VRB, which is specifically trained for kitchen tasks. Further, in its current form, the distilled visual grasp affordance does not depend on the orientation of the candidate grasp, limiting Grasp-Splat from fully harnessing grasp affordances in identifying better grasp configurations in $\mathrm{SE}(3)$. Future work will consider methods for extending the computation of grasp affordance to $\mathrm{SE}(3)$. Moreover, future work will seek to integrate online Gaussian Splatting into ASK-Splat, eliminating the need for a brief scanning phase of the scene prior to training. Additionally, we will seek to integrate sensor feedback into SEE-Splat to enable closed-loop, real-time scene editing, improving the accuracy of the scene representation. In addition, by closing the scene-editing loop with sensor feedback, we can ensure that our $3$D scene representation updates dynamically according to unexpected real-world events. For example, if the pan slips out of the robot gripper, SEE-Splat will be able to promptly reflect this change and modify the plans of the robot accordingly.

### v3

@@ B-A Semantic Localization via ASK-Splat
Given a natural-language query specifying an object of interest, SEE-Splat leverages ASK-Splat to identify semantically relevant Gaussians, as discussed in Section A-A. In the main paper (cf. Fig. 3), we show the localization of an electric stove, saucepan, and a fruit in a real-world Cooking scene. To improve the localization accuracy, the text prompt can include the geometric and visual properties of the object, such as its color, in addition to its semantic class. At this stage, SEE-Splat generates a semantic similarity map, from which relevant Gaussians are extracted, given a threshold on the semantic score.
B-B Masking the Gaussians in SEE-Splat
@@
Given a semantic similarity map of the scene, SEE-Splat generates a mask identifying the Gaussians relevant to the specified object. This procedure begins with thresholding the semantic scores of each Gaussian to remove dissimilar Gaussians from the set of relevant Gaussians, which creates a sparse point cloud of the relevant Gaussians, constructed from the means of these Gaussians. However, photo-realistic rendering of Gaussian environments require denser point clouds. Consequently, SEE-Splat lifts the features of the point cloud from the $3$D Euclidean space to a $7$D feature space, by augmenting each point in the point cloud with its RGB color and semantic score. Subsequently, SEE-Splat identifies all neighboring points in the scene within a specified distance of the point cloud in the $7$D feature space using an efficient KD-tree query. SEE-Splat incorporates these points into the point cloud to create a denser point cloud, comprising of all semantically-relevant Gaussians, while removing outliers from the set of points. In the main paper (cf. Fig. 3), we show the Gaussians extracted by SEE-Splat, as a point cloud with well-defined geometry, given a natural-language query for each object.
B-C Editing the Gaussians in SEE-Splat
@@
Leveraging the Gaussian primitives in ASK-Splat, SEE-Splat enables real-time scene-editing by inserting new Gaussians into the scene, removing Gaussians, and modifying the properties of the Gaussians, to reflect (or simulate) changes in the real-world. SEE-Splat supports seamless insertion and removal of Gaussians by introducing or deleting the relevant Gaussians from the set of Gaussians representing the scene, respectively. In addition, SEE-Splat supports both rigid and non-rigid transformation of the Gaussians, enabling simulated motion of the Gaussians, as well as changes to the shape of the Gaussians via non-isometric scaling. Specifically, given a function specifying the transformation ${\xi:\mathcal{G}_{s}\mapsto\mathcal{G}_{s}}$ (where $\mathcal{G}_{s}$ represents the space of the Gaussian primitives), SEE-Splat updates the spatial attributes of the relevant Gaussians by applying $\xi$ to these Gaussians. In the case of rigid transformations, ${\xi}$ can be described by an $\mathrm{SE}(3)$ transformation matrix, specifying rotation and translation of the Gaussians. We can render the edited scene to generate photo-realistic visualizations. Although, we do not consider physics-based simulations in this work, we note that physics can be incorporated into SEE-Splat to achieve realism. We expound on this point in or discussion on the limitations of SEE-Splat.
Deletion and transformation of the Gaussians introduces artifacts into the scene representation, degrading its photo-realistic qualities. To address this challenge, SEE-Splat enables $3$D Gaussian infilling by introducing new Gaussians with similar attributes in regions with missing geometry, which we illustrate in Appendix I. Figure 3 provides an illustration of such artifacts (e.g., the hole in the table), when the scene is edited to visualize the effects of moving the saucepan to the electric stove. Through $3$D Gaussian infilling, SEE-Splat generates a photorealistic rendering of the edited scene, eliminating these artifacts.
@@
Appendix C Grasping and Manipulation with Splat-MOVER
We present Grasp-Splat and discuss its application to multi-stage robotic manipulation via Splat-MOVER.
C-A Grasp-Splat for Grasp Proposal
@@
We note that the grasps generated by GraspNet are not always ideal. For example, the grasps generated by GraspNet in Figure 2 are either infeasible or challenging to execute. As a result, Grasp-Splat ranks the grasps proposed by GraspNet based on the grasp scores obtained from ASK-Splat. By leveraging the affordance score associated with each grasp pose, Grasp-Splat identifies grasp configurations that are more likely to succeed, depicted in Figure 2.
C-B Multi-Stage Robotic Manipulation
For multi-stage robotic manipulation, we begin by decomposing the manipulation task into stages. Our approach supports the specification of the stages comprising the task by a human or by a large language model (LLM). In the case where the natural-language description of the task does not specify the stages involved in the task, we query an LLM for the stages required to complete the manipulation task. For each stage in the manipulation task, we utilize ASK-Splat, SEE-Splat, and Grasp-Splat to identify the relevant object and generate candidate grasp poses. Likewise, we query ASK-Splat for the target location for placing the object. We evaluate the feasibility of each candidate grasp using an off-the-shelf motion planner for the robotic manipulator, inputting the point cloud of the scene, extracted from ASK-Splat and SEE-Splat, into the motion planner, which the motion planner uses for collision detection during motion planning. We execute the top candidate grasps, moving on to the next if the robot motion planner fails to compute a solution to execute the selected grasp.
@@ C-B Multi-Stage Robotic Manipulation
Appendix D Evaluations
We present additional experimental results of ASK-Splat, SEE-Splat, and Splat-MOVER in open-vocabulary, multi-stage robotic manipulation problems, including a discussion of the experimental setup.
D-A Experimental Setup
D-A1 ASK-Splat
@@ D-A3 Splat-MOVER
Specifically, we evaluate Splat-MOVER in four multi-stage manipulation tasks across three scenes: the Kitchen, Cleaning, and Workshop scenes. In the Kitchen scene, we consider a Cooking task where the robot is asked to place a saucepan on an electric burner (Stage 1) and subsequently place a fruit inside the saucepan (Stage 2). Further, in the Kitchen scene, we consider a Chopping task where the robot is asked to place a knife on a chopping board (Stage 1) and subsequently place a fruit next to the knife (Stage 2). We consider a Cleaning task (in the Cleaning scene), where the robot is asked to place a cleaning spray in a bin (Stage 1) and subsequently place a sponge next to the cleaning spray (Stage 2). Lastly, in the Workshop scene, the robot is asked to place a power drill on a work mat (Stage 1) and subsequently place a wooden block next to the drill (Stage 2), which we refer to as the Workshop task.
D-A4 Hardware Experiments
@@ D-B ASK-Splat Representation
D-C Splat-MOVER for Multi-Stage Robotic Manipulation
@@
We compare Splat-MOVER to prior work LERF-TOGO [5] and F3RM [6] in four tasks: the Cooking task, Chopping task, Cleaning task, and Workshop task, described in Section D-A3. Figure 7 shows a few candidate grasps proposed by GraspNet, F3RM, LERF-TOGO, and Grasp-Splat for each of these objects. GraspNet does not consider the semantic features of the object in generating candidate grasps; as a result, the proposed grasps are not localized in regions where a human might grasp the object, unlike the candidate grasps proposed by F3RM, LERF-TOGO, and Grasp-Splat, which generate grasps closer to the handle of the respective objects. For example, the proposed grasps lie relatively close to the handle of the saucepan and the knife. F3RM and LERF-TOGO generate candidate grasp conditioned on a text prompt identifying the region to grasp the object (such as its handle) provided by a human operator or an LLM (in LERF-TOGO) or from a dataset of human demonstrations (in F3RM). In contrast, Grasp-Splat does not require any external guidance to generate candidate grasps of similar quality, harnessing the grasp affordances provided by ASK-Splat. We summarize the capabilities of each of these methods in Table II.
@@
TABLE II: Representation Capabilities of LERF-TOGO [5], F3RM [6], and Splat-MOVER
Capabilities Semantic Knowledge Affordance Knowledge Scene-Editing
LERF-TOGO [5]
F3RM [6]
Splat-MOVER (ours)
In addition, we evaluate the pick-and-place success rate of all the methods in the Cooking task, Chopping task, Cleaning task, and Workshop task, where the place success rate is conditioned on the number of successful trials in picking the object. Table I provides the pick-and-place success rates in the Chopping task. Splat-MOVER achieves the highest pick success rate ($85\%$) in Stage $1$ of the task. Although F3RM achieves the highest place success rate, it achieves a much lower pick success rate compared to Splat-MOVER. In addition, in the Cleaning and Workshop tasks, Splat-MOVER achieves the highest success rates in the first stage of each task, and further achieves relatively high success rates in the second stage of each task, shown in Table I. LERF-TOGO achieves a perfect success rate in picking up the power drill in the first stage of the Workshop task. Since LERF-TOGO and F3RM are not amenable to multi-stage manipulation tasks, we cannot evaluate the success rate of these methods for the entire manipulation task. In contrast, Splat-MOVER enables multi-stage robotic manipulation, achieving a task success rate of $40\%$, $65\%$, $70\%$, and $80\%$ in the Cooking, Chopping, Cleaning, and Workshop tasks, respectively. We note that the Cooking task is the most challenging task, compared to the other tasks, given the little margin of error tolerated in placing the saucepan on the electric burner.